OpenStudy (loser66):

Given \(a_n = 2^n + 5*3^n\) for n =0,1,2,.... Prove that \(a_n = 5a_{n-1} -6a_{n-2} \) for \(n\geq 2\) Please, help

OpenStudy (anonymous):

i have no idea what happens if you try it directly?

OpenStudy (loser66):

What do you mean? I try induction, but I don't get the correct solution

OpenStudy (anonymous):

i mean just write out what \[5a_{n-1} -6a_{n-2}\] is

OpenStudy (anonymous):

it is not an induction proof i don't think

OpenStudy (anonymous):

in other words, write out \[5a_{n-1} -6a_{n-2}=5\left(2^{n-1}+5\times 3^{n-1}\right)-6\left(2^{n-2}+5\times 3^{n-2}\right)\]

OpenStudy (anonymous):

this is my best guess then muck around with the exponents

OpenStudy (loser66):

Ok, let me try .

OpenStudy (anonymous):

or as we say in elementary algebra, multiply out using the distributive property, combine like terms i think it should work

ganeshie8 (ganeshie8):

To simplify the algebra, you may use this trick : The given recurrence relation is linear. So, if \({r_1}^n\) and \({r_2}^n\) are two solutions, then by linearity it follows that all the linear combinations of those solutions are also solutions.

OpenStudy (loser66):

Can I use induction? But it doesn't work though. :(

OpenStudy (empty):

\[a_n = 5a_{n-1}-6a_{n-2}\] Plug 'em in: \[2^n+5*3^n = 5(2^{n-1}+5*3^{n-1}) - 6(2^{n-1}+5*3^{n-1})\]Rearrange the right hand side: \[2^n+5*3^n = \left( \frac{5}{2} - \frac{6}{4} \right) 2^n + \left( \frac{25}{3} - \frac{30}{9} \right) 3^n\]

ganeshie8 (ganeshie8):

Induction works nicely tooo

OpenStudy (loser66):

I got it. Thank you so much.

OpenStudy (loser66):

@Empty

OpenStudy (loser66):

@ganeshie8 Can you show me induction? Please

OpenStudy (empty):

Whoops, when I copied and pasted I just realized I didn't put in the -2 in the exponent I left it as -1 oh well cool glad it helped

ganeshie8 (ganeshie8):

By pluggin in, or by induction, you're just doing more of old highschool algebra.... you're not learning anything new...

OpenStudy (loser66):

I know and this is not my problem. I got mad when I can't find out the mistake I have on solving it for...... a high school student. But now I got what I missed. :)

ganeshie8 (ganeshie8):

Base case : \(a_0=6,a_1=17,a_2=49\) and \(49 = 5*17-6*6\) \(\checkmark\) Induction hypothesis : Assume \(a_n = 5a_{n-1}-6a_{n-2}\) is satisfied for the terms \(n=k,k-1\). Induction step : \(a_{k+1} = 2^{k+1}+5*3^{k+1}\\~\\ =2*2^k + 15*3^k\\~\\ =(5-3)*2^k+(25-10)*3^k\\~\\ =5*2^k+25*3^k- (3*2^k+10*3^k)\\~\\ =5(2^k+5*3^k) - 6(2^{k-1}+5*3^{k-1})\\~\\ =5a_k - 6a_{k-1}\\~\\ \blacksquare \)

OpenStudy (loser66):

Thank you very much.